OpenStudy (anonymous):

What is the slope of the graph of 2x – 6y = 9?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (lncognlto):

Haven't we gone over this...?

OpenStudy (skullpatrol):

$$\Huge 2x – 6y = 9$$ $$ \Huge 2x – 6y +6y= 9+6y$$ $$ \Huge 2x = 9+6y$$ $$ \Huge 2x-9 = 9+6y-9$$ $$ \Huge 2x-9 = 6y$$ $$ \Huge \frac{1}{6} ( 2x-9) = \frac{1}{6} 6y$$ $$ \Huge \frac{1}{3} x- \frac{3}{2} = y $$

OpenStudy (anonymous):

im sorry i got it by the last one but very time they ask somting diffreneet i just dont get it anyomore :(

OpenStudy (skullpatrol):

What don't you follow above?

OpenStudy (anonymous):

ok, tahnsk for trying

OpenStudy (anonymous):

ould it be -1 1/6?

OpenStudy (skullpatrol):

The answer is 1/3

OpenStudy (anonymous):

hmmm oh ok beacuse of the last step of the sketch you made me right? :)

OpenStudy (skullpatrol):

yes m = 1/3 in y = mx + b